Let the Utility Function be U = X3/4Y1/4 where PX = $1; PY = $2; I...

  1. Let the Utility Function be

U = X3/4Y1/4 where PX = $1; PY = $2; I = $15

Solve for Marshallian Demand (0.5 Points)

Solve for Hicksian Demand (0.5 Points)

What level of Utility is achieved at these quantities, prices and income? (0.5 Points)
Answer #1

classmate Date Page where Px = $1 pv = $2 I = $15 solue for Marshallian Demand. Budget line equation x +2y = 15 - =) slope ofJASMA -pue in Budget line equalian x+2y = 15 by+2y = 15 by = 15 = 15 X = 45 Marshallian Demand function for each google CX (PJassmate Date Page Solve for Hicksian Demand set Lagrange condition L = Px & Pay + x[uox L = x + 2y + x[ū - x zy & T FOC öx =RASSMALL | 6 - कप 41614 NP UGHicksian Demand function T c) achieved at utility as is and income. these quantities, prices, = (1.25) 2.4 (1083) = (6.1427)